### v6

@@ Abstract.
Proofs in proof assistants like Coq can be brittle, breaking easily in response to changes. To address this, recent work introduced an algorithm and tool in Coq to automatically repair broken proofs in response to changes that correspond to type equivalences. However, many changes remained out of the scope of this algorithm and tool—especially changes in underlying behavior. We extend this proof repair algorithm so that it can express certain changes in behavior that were previously out of scope. We focus in particular on equivalences between quotient types—types equipped with a relation that describes what it means for any two elements of that type to be equal. Quotient type equivalences can be used to express interesting changes in representations of mathematical structures, as well as changes in the underlying implementations of data structures.
We extend this algorithm and tool to support quotient type equivalences in Coq. Notably, since Coq lacks quotient types entirely, our extensions use Coq’s setoid machinery to represent quotients externally. Specifically, (1) our extension to the algorithm supports new changes corresponding to setoids, and (2) our extension to the tool supports this new class of changes and further automates away some of the new proof obligations. We demonstrate our extensions on proof repair case studies for previously unsupported changes. We also perform manual proof repair in Cubical Agda, a language with a univalent metatheory, which allows us to construct the first ever internal proofs of correctness for proof repair.
Proof Repair, Cubical Agda, Coq, Quotient Types
@@ 1. Introduction
Writing formal proofs in proof assistants like Coq, Agda, Lean, and Isabelle/HOL is a time-intensive task. Even once written, proofs may break in the face of minor changes in the datatypes, programs, and specifications they are about. User study data suggests that this process of writing and rewriting proofs is ubiquitous during proof development (Ringer et al., 2020), and that it can be challenging to deal with even for experts.
@@
Prior work introduced Pumpkin Pi, a Coq plugin for proof repair across changes in datatypes that can be described by type equivalences (Ringer et al., 2021). In this work, we extend proof repair to support equivalences between quotient types (Section 2). Recent work in Cubical Agda showed that certain relations describing changes in behavior can be adjusted to equivalences between quotient types (Angiuli et al., 2021). One specific example presented in that work uses two of the most basic queue representations. The first, one list queues, enqueue elements to the front of the list and dequeue elements from the back of the list. The second, two list queues, use a pair of lists. Elements are enqueued onto the front of the first list. When dequeuing, if the second list is empty, the first list is reversed onto the second, and then in any case the front element of the second list is removed. These types are not equivalent in any natural way, because multiple two list queues correspond to a single one list queue: for instance, the one list queue l corresponds to both (l, []) and ([], rev l). Angiuli et al. quotient two list queues so that, in the quotient type, (l, []) and ([], rev l) are in the same equivalence class, which makes the types equivalent. They are then able to perform proof transfer across that equivalence.
@@
We wish to use this same approach to implement proof repair instead of proof transfer. Unlike Cubical Agda, Coq lacks quotient types entirely, so one cannot use the original Pumpkin Pi transformation as-is to support this class of changes. To handle this, we represent quotient types externally using Coq’s setoid machinery, and we represent quotient type equivalences as setoid equivalences (Section 2). We then extend the proof transformation to support the newly generated equality proof obligations (Section 3).
@@
By extending Pumpkin Pi to quotient equivalences, users gain both expressive power and efficiency. Prior to this work, Pumpkin Pi could capture a similar approach using sigma types instead of setoids. To do this, one would choose a canonical element for each equivalence class in the quotient, and then use the subtype of these elements instead of a quotient or setoid. However, that approach has severe drawbacks. In the two list queue example above, when using both quotient and setoids representations, the multiple representatives of each class allow for an amortized constant time dequeue operation. The subtype representation instead only allows for a linear time dequeue operation.
@@
We implement the extended algorithm by way of an extension to the implementation of Pumpkin Pi (Section 4). Our implementation includes new automation, both to support repair across this class of changes and to automate away some of the newly generated proof obligations corresponding to this class of changes. We demonstrate our extended implementation on three case studies that cannot be handled by prior proof repair work—two that are mathematical in nature and one that deals with changes in behavior (Section 5). Finally, we define correctness for proof repair and construct internal proofs of correct repair in Cubical Agda (Section 6). Our contributions are:
@@
Our code is available in supplementary material.  [^1] https://github.com/uwplse/pumpkin-pi/tree/quotient-type-equivalences
@@ 2. Quotients and Setoids
Our goal is to extend Pumpkin Pi to perform repair between types equivalent up to quotients. To begin this task, we define quotient types:
@@
A quotient type A / eqA is defined by a type A and an equivalence relation eqA : A $\to$ A $\to$ Prop. A / eqA has one constructor, [ $\cdot$ ] : A $\to$ A / eqA, and its elements are called the equivalence classes of the elements of A. A / eqA has the additional property that, for a, b : A, [ a ] = [ b ] if eqA a b is inhabited. Eliminating an element q : A / eqA yields the underlying a : A from which it was constructed. However, the user must prove that the result of the computation is equal for any a, b : A with [ a ] = [ b ].
To give a concrete example, consider the quotient of $\mathbb{N}$ by the equivalence relation $\mathbb{N}$, where $\mathbb{N}$ n m is inhabited if n and m have the same parity. The resulting type, denoted by mathematicians as $\mathbb{N}/2$, has two elements, [ 0 ] and [ 1 ]; any other application of [ $\cdot$ ] produces a term equal to one of these. For $\mathbb{N}/2$, eliminating [ 0 ] and eliminating [ 2 ] must provably result in equal terms.
Quotient types can be challenging to implement within a language because of the equality property mentioned above. While some languages have native quotient types, many type systems, including Coq, do not support this property without the use of axioms. To circumvent this issue without introducing any axioms, we use setoids.
@@
Notice that setoids have no special constructors, eliminators, or equality properties. An element of A is said to be an element of the setoid, and users of the setoid should compare elements using eqA instead of native equality. Because of this, rewriting is more difficult in setoids than in quotient types. Rewriting by an equality in a quotient type can be done using the equality eliminator, but rewriting cannot generally be done for arbitrary equivalence relations. To compensate for this, we define the notion of a proper function.
@@
For setoids (A, eqA) and (B, eqB), a function f : A $\to$ B is proper if, for all a1, a2 : A, eqA a1 a2 implies that eqB (f a1) (f a2).
Using our example, $\mathbb{N}/2$, f : $\mathbb{N}$to $\mathbb{N}$ is proper if inputs to f of the same parity produce outputs of the same parity. The successor function would be one example of such a proper function. When a term is composed of proper functions, it becomes possible to construct proofs for rewriting by equivalence relations in much the same way we can rewrite by equality. Coq provides automation in its standard library to support constructing these proofs. By using setoids in place of quotients, equivalence relations in place of equality, and proper functions in place of arbitrary functions, we mimic the functionality of quotients using setoids.
Our paper deals with repair across quotient type equivalences. To begin, we define type isomorphisms:
@@
Types A and B are isomorphic if there exist functions f : A $\to$ B and g : B $\to$ A such that
@@
  • $\forall$ (b : B), f (g b) = b
Type equivalences are type isomorphisms with an additional adjoint property, which can be derived from any isomorphism. A quotient type equivalence is simply an equivalence between two quotient types. When using setoids to represent quotients, we must instead define a notion of a setoid equivalence:
@@
Two setoids (A, eqA) and (B, eqB) are equivalent if there exist functions f : A $\to$ B and g : B $\to$ A satisfying the following properties:
@@
Two setoids are equivalent if the quotient types they represent would be isomorphic. One example of a nontrivial setoid equivalence between the setoids ($\mathbb{N}$, $\mathbb{N}$) and (Bool, =) is given by the function isEven : $\mathbb{N}$to Bool, sending even numbers to true and odd numbers to false. This function has inverse g : Bool $\to \mathbb{N}$ mapping true to 0 and false to 1. f is proper, since isEven respects parity, and g is automatically proper because Bool uses equality as its equivalence relation. The other conditions, isEven (g b) = b and $\mathbb{N}$ (g (isEven n)) n, are both satisfied, and thus this indeed forms a setoid equivalence.
@@ 3. Approach: Proof Term Transformation
To reiterate, our goal is to perform proof repair across quotient type equivalences. Because we are working in Coq, which does not have quotient types, we represent quotients using setoids. Thus, given two types A and B which may either be types using standard equality or setoids using some equivalence relation, we wish to turn terms defined over A into terms defined over B which no longer have any references to A.
@@
Figure 1. The naturals represented in Coq, taken from the Pumpkin Pi paper (Ringer et al., 2021), in unary (left) and binary (right). In positive, xH is one, xO is appending a 0 to the right side of the binary representation, and xI is appending a 1 instead. Then, N is either 0 or a positive binary number.
@@ 3.1. Pumpkin Pi’s Transformation
The Pumpkin Pi transformation that we extend operates over terms in the the type theory of Coq, the Calculus of Inductive Constructions (CICω(Coquand and Paulin-Mohring, 1990). CICω extends the Calculus of Constructions (Coquand and Huet, 1988) with inductive types. The grammar for CICω is in Figure 2; the type rules are standard and omitted.
Pumpkin Pi implements proof repair over terms in CICω by transforming proof terms implemented over an old type A to instead be implemented over a new version of that type B. The key insight behind this transformation is that, by Lambek’s theorem, any equivalence between types A and B can be decomposed into separate components that talk only about A and only about B (Ringer, 2021). Functions and proofs can be unified with applications of these components, reducing repair to a simple proof term transformation replacing components that talk about A with their counterparts over B (Ringer et al., 2021).
@@
for types on both sides of the equivalence. [^2] While Pumpkin Pi has an $\eta$ term, meant to deal with $\eta$-expansions of constructors applied to eliminators, it appears to us that this term is not necessary. The example of nontrivial $\eta$ given in the Pumpkin Pi paper can be rewritten to instead have nontrivial $\iota$ and trivial $\eta$, and we are not aware of any cases where this cannot be done. Further, while the type of $\iota$ is always clearly defined, what $\eta$ should be is unclear for many examples. Thus, for the purposes of this paper, $\eta$ will always be trivial and we will ignore it. DepConstr and DepElim are, respectively, constructors and eliminators for each type, termed dependent constructors and dependent eliminators by Pumpkin Pi. The constructors must generate the elements of the type, and the eliminator must specify how to consume an element produced by the constructors. These constructors and eliminators must take the same shape, even if A and B themselves have different shapes.
For the type of unary naturals in Figure 1, we can choose depConstr for nat to be O and S, the constructors for nat. However, to repair to the binary naturals, we need to provide dependent constructors for N. These constructors must correspond across the equivalence, so we arrive at constructors with these types:
@@
These dependent constructors do not share the type signatures of N’s constructors, but rather are user-defined functions corresponding to nat’s constructors. Likewise, we can choose the eliminator for nat, nat_rect, as the depElim for nat, and the dependent eliminator for N then takes the shape of nat’s eliminator:
@@
We give a concrete example of repairing addition in Figure 3.
@@
The fact that both dependent eliminators must have the same shape even when the underlying types do not is exactly why we need the remaining element of the configuration: $\iota$. This gives the $\iota$-reduction rules, which specify how to reduce an application of a dependent eliminator to a dependent constructor. When the shape is of the underlying type is the same as the shape of the configuration components, as is true for for nat, this will be definitional—the proof assistant will handle it automatically. However, if the inductive structure has changed, as it has for N, this will be a propositional equality.
@@
flushleft $\Gamma$ $\vdash$ $t$Uparrow $t^{\prime}$

### v7

@@ Abstract.
Proofs in proof assistants like Rocq can be brittle, breaking easily in response to changes. To address this, recent work introduced an algorithm and tool in Rocq to automatically repair broken proofs in response to changes that correspond to type equivalences. However, many changes remained out of the scope of this algorithm and tool—especially changes in underlying behavior. We extend this proof repair algorithm so that it can express certain changes in behavior that were previously out of scope. We focus in particular on equivalences between quotient types—types equipped with a relation that describes what it means for any two elements of that type to be equal. Quotient type equivalences can be used to express interesting changes in representations of mathematical structures, as well as changes in the implementations of data structures.
We extend this algorithm and tool to support quotient type equivalences in Rocq. Notably, since Rocq lacks quotient types entirely, our extensions use Rocq’s setoid machinery in place of quotients. Specifically, (1) our extension to the algorithm supports new changes corresponding to setoids, and (2) our extension to the tool supports this new class of changes and further automates away some of the new proof obligations. We demonstrate our extensions on proof repair case studies for previously unsupported changes. We also perform manual proof repair in Cubical Agda, a language with a univalent metatheory, which allows us to construct the first ever internal proofs of correctness for proof repair.
Proof Repair, Cubical Agda, Rocq, Quotient Types
copyright: ccdoi: 10.1145/3763164journalyear: 2025journal: PACMPLjournalvolume: 9journalnumber: OOPSLA2article: 386publicationmonth: 10ccs: Software and its engineering Software evolutionccs: Theory of computation Type theory
@@ 1. Introduction
Writing formal proofs in proof assistants like Rocq [^1] Formerly known as Coq., Agda, Lean, and Isabelle/HOL is a time-intensive task. Even once written, proofs may break in the face of minor changes in the datatypes, programs, and specifications they are about. User study data suggests that this process of writing and rewriting proofs is ubiquitous during proof development (Ringer et al., 2020), and that it can be challenging to deal with even for experts.
@@
Prior work introduced Pumpkin Pi, a Rocq plugin that performs proof repair across changes in datatypes that can be described by type equivalences (Ringer et al., 2021). In this work, we extend proof repair to support equivalences between quotient types (Section 2). Recent work by Angiuli et al. (2021) in Cubical Agda showed that certain relations describing changes in behavior can be adjusted to equivalences between higher inductive types. One specific example presented in that work uses two queue representations. The first, one list queues, enqueue elements to the front of the list and dequeue elements from the back of the list. The second, two list queues, use a pair of lists. Elements are enqueued onto the front of the first list. When dequeuing, if the second list is empty, the first list is reversed onto the second, and then in any case the front element of the second list is removed. These types are not equivalent in any natural way, because multiple two list queues correspond to a single one list queue: for instance, the one list queue l corresponds to both (l, []) and ([], rev l). Angiuli et al. use a higher inductive type to construct the quotient type of two list queues so that, in the resulting quotient type, (l, []) and ([], rev l) are in the same equivalence class, which makes the types equivalent. They are then able to perform proof transfer across that equivalence.
@@
We wish to use this same approach to implement proof repair instead of proof transfer. Unlike Cubical Agda, Rocq lacks quotient types entirely, so one cannot use the original Pumpkin Pi transformation as-is to support this class of changes. To handle this, we replace quotient types by using Rocq’s setoid machinery, and we replace quotient type equivalences with setoid equivalences (Section 2). We then extend the proof transformation to support the newly generated equality proof obligations (Section 3).
@@
By extending Pumpkin Pi to setoids, users gain both expressive power and efficiency. Prior to this work, Pumpkin Pi could capture a similar approach using sigma types instead of setoids. To do this, one would choose a canonical element for each equivalence class in the quotient, and then use the subtype of these elements instead of a quotient or setoid. However, that approach has severe drawbacks. In the two list queue example above, when using both quotient and setoids representations, the multiple representatives of each class allow for an amortized constant time dequeue operation. The subtype representation instead only allows for a linear time dequeue operation.
@@
We implement this algorithm by way of an extension to the implementation of Pumpkin Pi (Section 4). Our implementation includes new automation, both to support repair across this class of changes and to automate away some of the newly generated proof obligations corresponding to this class of changes. We demonstrate our extended implementation on three case studies that cannot be handled by prior proof repair work—two that are mathematical in nature and one that deals with changes in behavior (Section 5). Finally, we define correctness for proof repair and construct internal proofs of correct repair in Cubical Agda (Section 6). Our contributions are:
@@
Our code is available on Github, and we have packaged the extension, our case studies, and all the tools required to run them into an artifact (Viola et al., 2025) for this paper. [^2] Throughout the paper, links to specific files in the Github source have been provided where relevant. These links are given as circled numbers, like 1.
@@ 2. Quotients and Setoids
Our goal is to extend Pumpkin Pi to perform repair between types equivalent up to quotients. To begin this task, we define quotient types (Hofmann, 1995):
@@
A quotient type A / eqA is defined by a type A and an equivalence relation eqA : A $\to$ A $\to$ Prop. A / eqA has one constructor, [ $\cdot$ ] : A $\to$ A / eqA, and its elements are called the equivalence classes of the elements of A. A / eqA has the additional property that, for a, b : A, [ a ] = [ b ] if eqA a b is inhabited. Eliminating an element q : A / eqA yields the underlying a : A from which it was constructed. However, the user must prove that the result of the computation that is being defined is equal for any a, b : A with [ a ] = [ b ].
[…]
Quotient constructions have historically been challenging to implement directly in constructive logic. Because of the equality property mentioned above, many type systems, including Rocq, do not support quotient types without the use of axioms. Adding quotient types can have nonobvious consequences as well. For instance, extending intensional Martin-Löf type theory with effective quotients and uniqueness of identity proofs is sufficient to prove the law of the excluded middle for types in the first universe if the type theory has at least two universes (Maietti, 1999). A common alternative approach, introduced by Bishop (Bishop, 1967) for constructive analysis and popularized in type theory by Hofmann (1994, 1995), is to use setoids.
@@
Alternative definitions exist; see Barthe et al. (2003) for a survey of these definitions. Setoids are a natural alternative when actually constructing a quotient, and in turn its equivalence classes, is unfeasible. The Rocq standard library provides both definitions for setoids and substantial automation for their use. Quotients, on the other hand, see no such support. Their implementation would require the addition of axioms, which block computation and are generally avoided by Rocq users when possible. Thus, in Rocq, using setoids in place of quotients is the norm (Team, 2023; Ringer et al., 2019). Furthermore, using axioms in a Rocq plugin forces users of that tool to adopt the axiom, so using axioms in plugins is doubly discouraged.
Notice that setoids have no special constructors, eliminators, or equality properties. An element of A is said to be an element of the setoid, and users of the setoid should compare elements using eqA instead of native equality. Because of this, rewriting is more difficult in setoids than in quotient types. Rewriting by an equality in a quotient type can be done using the equality eliminator, but rewriting cannot generally be done for arbitrary equivalence relations. To compensate for this, we need the notion of a proper function.
@@
For setoids (A, eqA) and (B, eqB), a function f : A $\to$ B is proper if, for all a1, a2 : A, eqA a1 a2 implies that eqB (f a1) (f a2). (Sozeau, 2023; Team, 2019)
Using our example, $\mathbb{N}/2$, f : $\mathbb{N}$ $\to$ $\mathbb{N}$ is proper if inputs to f of the same parity produce outputs of the same parity. The successor function would be one example of such a proper function. When a term is composed of proper functions, it becomes possible to construct proofs for rewriting by equivalence relations in much the same way we can rewrite by equality. Rocq provides automation in its standard library to support constructing these proofs. By using setoids in place of quotients, equivalence relations in place of equality, and proper functions in place of arbitrary functions, we mimic the functionality of quotients using setoids.
Our paper deals with repair across quotient type equivalences. To begin, we recall the definition of a type isomorphisms:
@@
An isomorphism from type A to type B is a function f : A $\to$ B such that there exists g : B $\to$ A with
@@
  • $\forall$ (b : B), f (g b) = b (Univalent Foundations Program, 2013)
Type equivalences are type isomorphisms with an additional adjoint property, which can be derived from any isomorphism. A quotient type equivalence is simply an equivalence between two quotient types. When using setoids to represent quotients, we must instead use a notion of a setoid equivalence:
@@
A setoid equivalence between (A, eqA) and (B, eqB) is a function f : A $\to$ B such that there exists g : B $\to$ A satisfying the following properties:
@@
[…]
@@ 3. Approach: Proof Term Transformation
To reiterate, our goal is to perform proof repair across quotient type equivalences. Because we are working in Rocq, which does not have native quotient types, we represent quotients using setoids. Thus, given two types A and B, each of which the user may choose to represent as either a type using standard equality or a setoid using some equivalence relation, we wish to turn terms defined over A into terms defined over B which no longer have any references to A.
@@
Figure 1. The natural numbers, as represented in Rocq’s standard library, in unary (left) and binary (right). The type positive represents all positive natural numbers. xH is one, xO n is appending a 0 to the right side of the binary representation of n, and xI n is appending a 1 to the right side of the binary representation of n. Then, N is either 0, as N0, or a positive binary number.
@@ 3.1. Pumpkin Pi’s Transformation
The Pumpkin Pi transformation that we extend operates over terms in the type theory of Rocq, the Calculus of Inductive Constructions (CICω(Coquand and Paulin-Mohring, 1990). CICω extends the Calculus of Constructions (Coquand and Huet, 1988) with inductive types. The grammar for CICω is in Figure 2; the type rules are standard and omitted.
Pumpkin Pi implements proof repair over terms in CICω by transforming proof terms implemented over an old type A to instead be implemented over a new version of that type B. The key insight behind this transformation is that any equivalence between types A and B can be decomposed into separate components that talk only about A and only about B (Ringer, 2021). Functions and proofs can be unified with applications of these components, reducing repair to a simple proof term transformation replacing components that talk about A with their counterparts over B (Ringer et al., 2021).
@@
for types on both sides of the equivalence. DepConstr and DepElim are, respectively, constructors and eliminators for each type, termed dependent constructors and dependent eliminators internally by Pumpkin Pi (even in cases where those constructors and eliminators might be non-dependently typed). The constructors must generate the elements of the type, and the eliminator must specify how to consume an element produced by the constructors. These constructors and eliminators must take the same shape, even if A and B themselves have different shapes. Usually, then, for at least one of the types A or B DepConstr and DepElim will be distinct from the usual constructors and eliminators for that type.
Like the Pumpkin Pi paper (Ringer et al., 2021), we will use repair between unary and binary natural numbers as an example. For the type of unary naturals in Figure 1, we can choose depConstr for nat to be O and S, the constructors for nat. However, to repair to the binary naturals, we need to provide dependent constructors for N. These constructors must correspond across the equivalence, so we must provide constructors with these types:
@@
These dependent constructors do not share the type signatures of N’s constructors, but rather are user-defined functions corresponding to nat’s constructors. Specifically, the successor constructor for N instead has type positive -> N. Likewise, we can choose the eliminator for nat, nat_rect, as the depElim for nat, and the dependent eliminator for N then must take the shape of nat’s eliminator:
@@
We give a concrete example of repairing addition in Figure 3   1.
@@
The fact that both dependent eliminators must have the same shape even when the underlying types do not is exactly why we need the remaining element of the configuration: $\iota$. This gives the $\iota$-reduction rules, which specify how to reduce an application of a dependent eliminator to a dependent constructor. When the shape is of the underlying type is the same as the shape of the configuration components, as is true for for nat, this will be definitional—the proof assistant will handle it automatically. However, if the inductive structure is different, as it is for N, this will be a propositional equality. As an example, the $\iota$-reduction rule for depConstrNSuc has the following type:
Definition iotaNSuc (P : N -> Type) (PO : P 0)
(PS : forall x : N, P x -> P (depConstrNSuc x))
(n : N) (Q : P (depConstrNSuc n) -> Type) :
Q (PS n (depElimN P PO PS n)) -> Q (depElimN P PO PS (depConstrNSuc n))
We may also use a version of this rule in proofs which runs in the reverse direction:
Definition iotaNSucRev (P : N -> Type) (PO : P 0)
(PS : forall x : N, P x -> P (depConstrNSuc x))
(n : N) (Q : P (depConstrNSuc n) -> Type) :
Q (depElimN P PO PS (depConstrNSuc n)) -> Q (PS n (depElimN P PO PS n))
The last term, $\eta$, represents reducing the $\eta$-expansion of constructors applied to eliminators. However, we have yet to find an example where a nontrivial $\eta$ is actually necessary. Thus, for the purposes of this paper, $\eta$ will always be trivial and we will ignore it.  [^3] The example of nontrivial $\eta$ given in the Pumpkin Pi paper can be rewritten to instead have nontrivial $\iota$ and trivial $\eta$   2, and we are not aware of any cases where this cannot be done. Further, while the type of $\iota$ is always clearly defined, what $\eta$ should be is unclear for many examples. For this reason, it’s possible that $\eta$ should be dropped entirely, and a configuration should simply include DepConstr, DepElim, and $\iota$.
@@
flushleft
$\Gamma$ $\vdash$ $t$ $\Uparrow$ $t^{\prime}$